A child standing on the floor of a balcony of a multi storey building, throws a ball straight upwards with an initial velocity of 10m/s. Assume the the floor is 25m high from the ground. Take g=9.8m/s/s.find the a.maximum height above the ground reached by the ball. B.velocity with which the ball strikes the ground
1
Expert's answer
2019-11-04T10:36:34-0500

A.


"H=h+\frac{v^2}{2g}"

"H=25+\frac{10^2}{2(9.8)}=30\ m"

B.


"V^2=v^2+2gh"

"V^2=10^2+2(9.8)(25)"

"V=24\frac{m}{s}"